OpenStudy (anonymous):

How many yards of material from a 24 yard length of cloth remain after 3 pieces, each 3 1/2 yards long, and 5 pieces each 2 1/4 yards long are removed?

OpenStudy (anonymous):

anyone know the answer

OpenStudy (gw2011):

The answer is 2.25 = 2 1/4. Here is how the answer was arrived at: (3)(3.5) + (5)(2.25) = 10.5 + 11.25 = 21.75 24 - 21.75 = 2.25 = 2 1/4